# Topolectrical circuits – recent experimental advances and developments

### v1

@@ III.2 Non-Hermitian skin effect
- III.2.6 Circuit setups for non-Hermitian sensing
@@ I Introduction
Low loss and component uncertainties: That said, many circuit component manufacturers provide a wide variety of high-quality components with minimal loss and deviations in component parameters. Additionally, for implementations that demand exceptionally low loss and deviation, it is feasible to preselect circuit elements to meet these stringent requirements due to their generally low cost.
Direct observability and measurability: Impedance measurements allow key band structure features, such as non-trivial topological zero modes, to be directly observed through straightforward profiling [45]. Such measurements can either measure the local transmissibility or the global, non site-specific band structure properties [128].
@@
Although some of these advantages may be even more prominent in certain platforms i.e. strong intrinsic optical non-linearity in non-linear photonics [140, 141, 142, 143, 144], electrical circuits remain a highly suitable medium for realizing a broad range of phenomena, as outlined above. In general, their versatility and practicality make them an excellent choice for many applications.
@@ II.1 Laplacian formalism for circuits
where and are the sub-lattice nodes in a unit cell , represent the admittances of the circuit node connections, and , with denoting the admittance of the grounding components. For example, in the SSH circuit with inter-cell and intra-cell capacitances and , and grounding inductance , the admittances are , , and , where and is the frequency of the driving signal. While these coupled equations also occur in the commonly known SSH model Hamiltonian, capturing the hoppings between sub-lattice nodes and within each unit cell , we emphasize its distinction to the Hamiltonian describing electrical circuits, as described below.
@@ II.2.1 Method of images for bounded circuit arrays
Importantly, can be extraordinarily large for certain special values of certain circuit sizes if there are circuit coupling elements that introduces opposite phases i.e. capacitors and inductors [Fig. 1e]. This is because having at least two different of opposite signs can allow the denominator to almost vanish at appropriate values of . Such dramatic system size dependence in the impedance is not seen in homogeneous electrical media, where the dependence on is smooth (logarithmic or power-law) [17, 161, 162, 163].
Behind the seemingly erratic occurrence of impedance peaks in Fig. 1e is a sophisticated fractal-like pattern of the resonance height in the parameter space of system size and , the dimensionless (normalized) frequency characterized by the LC components [Fig. 1f]. While the detailed pattern depends on the actual lattice model describing the circuit array, hierarchies of impedance peaks universally occur in a qualitatively similar manner. The intricacy of this pattern arises from the commensurability properties of the denominator in Eq. (10), in which some values of can cause it to be especially close to vanishing [Fig. 1g]. Notably, there exists some branches that are particularly robust and hence experimentally detectable [153], reminiscent of the more salient structures in the Hofstadter butterfly. Additionally, resonances may also arise due to protected topological zero modes, as in a Chern lattice.
@@ II.3.2 Through direct circuit resonance measurement
Constructing the circuit Laplacian in full requires impedance measurements in general, and that presents practical limitations in larger circuits. Recently, Franca et al. [128] developed a method to extract the Laplacian eigenenergies directly from impedance resonances using much fewer measurement configurations. As deduced from Eq. (9), any vanishing eigenvalue corresponds to a large impedance resonance. By tuning the system across a range of AC frequencies, various resonances can observed, as demonstrated with high accuracy in their Fibonacci circuit implementation [Fig. 1h]. This approach bypasses the need for constructing the circuit Laplacian, reducing the measurement complexity from measurements to a single impedance measurement at various (easily tunable) frequencies.
To elaborate, this approach involves examining the second derivative of the impedance profile with respect to frequency, coupled with a fourth-order Butterworth filter to mitigate noise [Fig. 1i]. Advanced data processing techniques, such as the matrix pencil method and utilizing the symmetric energy profiles of chiral symmetric systems, further enhance the precision in determining admittance eigenvalues. In a similar spirit as other studies that utilize machine learning to infer the spectrum from limited measurements (to be discussed in subsequent sections e.g., Ref. [169]), this method represents a significant advance which allows for the the measurement of the Laplacian spectrum from a single measurement [Fig. 1j].
@@ III.1.1 Exceptional points in parity-time symmetric setups
Parity-time (PT) symmetry can protect the reality of eigenvalues even if the operator i.e. circuit Laplacian or Hamiltonian is non-Hermitian [176, 177, 178, 179, 180, 181]. A typical PT-symmetric 2-level Hamiltonian is represented as
@@
The eigenvalues of are given by . When , the eigenvalues are real, indicating the PT-symmetric phase. Conversely, when , the eigenvalues become complex, signaling a transition to the broken PT-symmetric phase. At the critical point , not only do the eigenvalues coincide, but the eigenvectors also coalesce (become parallel to each other). Such non-Hermitian critical points are known as exceptional points (EPs) [182, 183, 184, 185, 186, 187, 188, 180], which are special because the eigenspace is defective i.e. not full rank. Further, due to the square-root singularity of the eigenbands at the EP, the spectrum becomes highly sensitive to perturbations, with a divergent gradient as the EP is approached. This sensitivity increases for higher-order exceptional points, where eigenvalues coincide and result in a singularity. That the system’s response to perturbations becomes increasingly pronounced is particularly advantageous for sensing applications [189, 190]. The sensitivity can be enhanced by incorporating multiple modes or alternative coupling mechanisms such as to enable higher-order EPs [188, 191, 192, 193, 194, 195]. Beyond isolated EPs, higher dimensional extensions such as exceptional lines [196] and exceptional rings [197] have also been observed, in analogy to nodal lines and rings [198, 199, 200, 201, 202, 203].
@@
Although an EP system becomes highly responsive at the EP, the accompanying defective eigenspace (where two or more eigenvectors coalesce) can also introduce undesirable susceptibility to noise [206], complicating the implementation of EP-based sensors [183]. However, appropriately designed non-linearity can mitigate that susceptibility while preserving the responsiveness, even leading to EP transitions [207, 208]. Bai et al. [173] designed a two-component circuit that juxtaposes the non-linearity and non-Hermitian coupling between two oscillators [Fig. 2d1]. Introducing non-linear gain results in higher-order non-linear EPs (NEPs), such as third-order NEPs as shown in Fig. 2d2. They effectively suppresses noise, even when the circuit is intentionally disturbed by a non-negligible external pulse [209]. With non-linearity, the Petermann factor can be prevented from diverging [Fig. 2d3]. Additionally, the signal-to-noise ratio reaches its lowest value at the NEP and increases as the system moves away from the NEP [Fig. 2d4].
PT symmetry-based implementations can also enhance system stability. In semiconductor technologies, fast-switching devices are widely used, and achieving both speed and stability is highly desirable [210, 211]. Yang et al. [174] demonstrated that transient PT symmetry, as implemented in their electric circuit [Fig. 2e1], can significantly reduce unwanted oscillations during fast switching. Noise-induced perturbations in the oscillations decay more rapidly at the EP compared to in the broken or unbroken PT phases [Fig. 2e2]. For step-like dynamic switching, transient PT symmetry effectively suppresses unwanted oscillations [red in Fig. 2e3], as observed from the diminishing peak around 36 MHz in the transient PT phase [blue in Fig. 2e4] in the frequency domain.
Mathematically, EPs can also trace out interesting mathematical structures in either parameter or real/momentum space. EPs underlie braiding in the complex eigenvalue plane [212, 213] and define knots in the band topology [214]. By designing non-Hermitian circuits involving multiple bands, Cao et al. [175] achieved higher-order EPs. The shape of knots or braids primarily depends on the winding paths around the EPs [215]. For example, a circuit with two nodes [Figs. 2f1 and f2] or three nodes [Fig. 2f3] can exhibit two-component or three-component knot topology, with double or triple twisting in the complex eigenenergy plane. Momentum-space EP knots are also described in Refs. [216, 217, 218]. In a larger parameter space, multi-band EPs can also describe mathematical singularities known as catastrophes, which has deep connections i.e. Mckay correspondence with diverse mathematical constructs such as the ADE classification of Lie algebras, finite subgroups of , platonic solids and the monodromy group of simple singularities [219, 220]. A recent circuit experiment has mapped out the swallowtail catastrophe in parameter space [221].
@@ III.1.2 Exceptional bound states as robust circuit resonances
The defectiveness of EPs causes the occupied band projector to be singular, leading to substantial non-locality in its two-point correlation functions. A prominent consequence is the appearance of special isolated eigenstates [Fig. 3a] of the real-space truncated two-point correlator matrix known as exceptional bound (EB) states [223, 222, 224, 225]. These EB states differ from topological modes, which are protected by non-trivial bulk topology, and non-Hermitian skin modes, which stem from non-trivial point gap band topology and are also boundary-localized in open-boundary systems [226, 227, 228, 229, 230, 231, 232, 233]. While most eigenvalues of the truncated two-point correlator matrix remain clustered around or , as in non-EB cases, a distinctive pair of EB states and additionally appears as long as the parent Hamiltonian is defective. The value of scales strongly with the system size, as detailed in Refs. [223, 222]. Since the spectrum of the truncated two-point correlator matrix corresponds to fermionic occupation probabilities, EB states also lead to negative entanglement entropy in the context of free fermions [223, 222, 234], with interesting ramifications in topological [225] or competitive non-Hermitian skin effect settings [235].
@@ III.2 Non-Hermitian skin effect
One of the most intensely studied non-Hermitian phenomenon is the the non-Hermitian skin effect (NHSE), where states accumulate against spatial inhomogeneities, impurities or boundaries [236, 237, 238, 239, 240, 241, 228, 229, 226, 242, 165, 243, 244, 245, 246, 247, 248, 249, 250, 251, 252, 253, 254, 255, 256, 257, 258, 259, 260, 261, 262, 263, 264, 265, 266, 267, 268, 269, 270, 271, 272, 273, 274, 275, 276, 277, 278, 279, 280, 281, 282, 283, 284, 285], challenging conventional condensed-matter insights into the correspondence between bulk and boundary properties.
In most cases, the NHSE occurs when the periodic boundary condition (PBC) spectrum encloses a finite area in the 2D complex spectral plane, which generally occurs when asymmetric couplings are present. Then, it can be shown that the open boundary condition (OBC) spectrum must assume the form of arcs or branches within the PBC spectral loop. Referring the reader to the many excellent references that elaborate on this [286, 227, 230, 287, 288, 233, 240, 289, 244, 290, 291], we briefly note that the OBC spectrum, and hence the NHSE, can be understood from the Bloch PBC Hamiltonian by complex-deforming the momentum i.e. introducing an imaginary flux [241, 292, 247, 291]:
where and represent the real and imaginary components of the momentum wavevector, respectively. The form of the complex deformation , which generically depend on , is crucial for understanding the nature of the so-called skin modes which exponentially localize against the system’s boundaries. For any Bloch wavevector with eigenvector , an appropriate complex momentum deformation would yield the OBC eigenequation
@@
Because is in general not in the Bloch (PBC) spectrum, is in general nonzero too. To determine values of which are actually allowed to exist in the OBC spectrum, we note that satisfying OBCs at both boundaries simultaneously generally require the exact OBC eigenstate to be a superposition of two different bulk solutions with equal decay lengths . In other words, the OBC eigenenergies are those values of for which Eq. (14) simultaneously admits two different solutions with the same value of , a condition famously known as the GBZ condition [238, 293, 294, 295, 291].
Note that the GBZ condition, as well as the existence of an unique that quantifies the skin accumulation, may no longer hold when more than one NHSE channel coexists in the system. A prominent example of which is the critical non-Hermitian skin effect (cNHSE) [296, 297, 298, 235, 299, 300, 224], where weakly coupling chains with different NHSE directions produces feedback loops that amplify certain states and lead to qualitatively different band structure descriptions at different system sizes. Appropriately designed cNHSE systems may even undergo an exceptional phase transition at a particular system size, with special scaling-induced EP defectiveness and lead to unique entanglement dips [224]. Fundamentally, the GBZ condition hinges on the assumption that the OBC eigenstate only comprise of the superposition of two bulk solutions; relaxing that generally leads to a “fragmented” GBZ in which more than one participates in the physics substantially [301, 302].
@@ III.2.1 Circuit demonstrations of one-dimensional NHSE phenomena
In a related vein, Shen et al. [304] conducted a circuit experiment involving multiple asymmetric hoppings in coupled 1D SSH chains [Fig. 4c1]. They extended this setup by connecting multiple 1D SSH chains to form a spherical geometry, paving the framework for studying the interplay between band topology and real-space lattice topology [166]. While localization behavior in various lattice geometries, including kagome and honeycomb lattices [309, 310], have been widely studied, in the spherical model [Fig. 4c2, c3], localization occurs only when a defect is introduced [pointed by red arrow in Fig. 4c4]. Additionally, Jiang et al. [311] demonstrated how circuits with multiple non-reciprocal couplings can give rise to corner modes in 2D circuits.

### v2

@@ III.2 Non-Hermitian skin effect
- III.3 Circuit applications for non-Hermitian sensing
@@ I Introduction
Precise control of component uncertainties and loss: That said, many circuit component manufacturers provide a wide variety of high-quality components with minimal deviations in component parameters and loss. Additionally, for implementations that demand exceptionally low loss and deviation, it is feasible to preselect circuit elements due to their generally low cost, as a variety of low-loss components can be readily found to meet these stringent requirements.
Direct observability and measurability: Impedance measurements allow key band structure features, such as non-trivial topological zero modes, to be directly observed through straightforward profiling [45]. Such measurements can either measure the local transmissibility or the global, non-site-specific band structure properties [128].
@@
Although some of these advantages may be even more prominent in certain platforms, i.e., strong intrinsic optical non-linearity in non-linear photonics [140, 141, 142, 143, 144], electrical circuits remain a highly suitable medium for realizing a broad range of phenomena, as outlined above. In general, their versatility and practicality make them an excellent choice for many applications.
@@ II.1 Laplacian formalism for circuits
where and are the sub-lattice nodes in a unit cell ; represent the admittances of the circuit node connections, and , with denoting the admittance of the grounding components. For example, in the SSH circuit with inter-cell and intra-cell capacitances and , and grounding inductance , the admittances are , , and , where and is the frequency of the driving signal. While these coupled equations also occur in the commonly known SSH model Hamiltonian, capturing the hoppings between sub-lattice nodes and within each unit cell , we emphasize its distinction to the Hamiltonian describing electrical circuits, as described below.
@@ II.2.1 Method of images for bounded circuit arrays
Importantly, can be extraordinarily large for certain some values of certain circuit sizes if there are circuit coupling elements that introduce opposite phases, i.e., capacitors and inductors [Fig. 1e]. This is because having at least two different of opposite signs can allow the denominator to almost vanish at appropriate values of . Such dramatic system size dependence in the impedance is not seen in homogeneous electrical media, where the dependence on is smooth (logarithmic or power-law) [17, 161, 162, 163].
Behind the seemingly erratic occurrence of impedance peaks in Fig. 1e is a sophisticated fractal-like pattern in the distribution of resonance heights within the parameter space of system size and —the dimensionless (normalized) frequency characterized by the LC components [Fig. 1f]. While the detailed pattern depends on the actual lattice model describing the circuit array, hierarchies of impedance peaks universally occur in a qualitatively similar manner. The intricacy of this pattern arises from the commensurability properties of the denominator in Eq. (10), in which some values of can cause it to be especially close to vanishing [Fig. 1g]. Notably, there exists some branches that are particularly robust and hence experimentally detectable [153], reminiscent of the more salient structures in the Hofstadter butterfly. Additionally, resonances may also arise due to protected topological zero modes, as in a Chern lattice [bottom diagram in Fig. 1f].
@@ II.3.2 Through direct circuit resonance measurement
Constructing the circuit Laplacian in full requires impedance measurements in general, and that presents practical limitations in larger circuits. Recently, Franca et al. [128] developed a method to extract the Laplacian eigenenergies directly from impedance resonances using much fewer measurement configurations. As deduced from Eq. (9), any vanishing eigenvalue corresponds to a large impedance resonance. By tuning the system across a range of AC frequencies, various resonances can be observed, as demonstrated with high accuracy in their Fibonacci circuit implementation [Fig. 1h]. This approach bypasses the need for constructing the circuit Laplacian, reducing the measurement complexity from measurements to a single impedance measurement at various (easily tunable) frequencies.
To elaborate, this approach involves examining the second derivative of the impedance profile with respect to frequency, coupled with a fourth-order Butterworth filter to mitigate noise [Fig. 1i]. Advanced data processing techniques, such as the matrix pencil method and utilizing the symmetric energy profiles of chiral symmetric systems, further enhance the precision in determining admittance eigenvalues. In a similar spirit as other studies that utilize machine learning to infer the spectrum from limited measurements (to be discussed in subsequent sections, e.g., Ref. [169]), this method represents a significant advance which allows for the the measurement of the Laplacian spectrum from a single measurement [Fig. 1j].
@@ III.1.1 Exceptional points in parity-time symmetric setups
Parity-time (PT) symmetry can protect the reality of eigenvalues even if the operator, i.e., circuit Laplacian or Hamiltonian is non-Hermitian [176, 177, 178, 179, 180, 181]. A typical PT-symmetric 2-level Hamiltonian is represented as
@@
The eigenvalues of are given by . When , the eigenvalues are real, indicating the PT-symmetric phase. Conversely, when , the eigenvalues become complex, signaling a transition to the broken PT-symmetric phase. At the critical point , not only do the eigenvalues coincide, but the eigenvectors also coalesce (become parallel to each other). Such non-Hermitian critical points are known as exceptional points (EPs) [182, 183, 184, 185, 186, 187, 188, 180], which are special because the eigenspace is defective, i.e., not full rank. Further, due to the square-root singularity of the eigenbands at the EP, the spectrum becomes highly sensitive to perturbations, with a divergent gradient as the EP is approached. This sensitivity increases for higher-order exceptional points, where eigenvalues coincide and result in a singularity. That the system’s response to perturbations becomes increasingly pronounced is particularly advantageous for sensing applications [189, 190]. The sensitivity can be enhanced by incorporating multiple modes or alternative coupling mechanisms to enable higher-order EPs [188, 191, 192, 193, 194, 195, 196]. Beyond isolated EPs, higher dimensional extensions such as exceptional lines [197] and exceptional rings [198] have also been observed, in analogy to nodal lines and rings [199, 200, 201, 202, 203, 204].
@@
Although an EP system becomes highly responsive at the EP, the accompanying defective eigenspace (where two or more eigenvectors coalesce) can also introduce undesirable susceptibility to noise [207], complicating the implementation of EP-based sensors [183]. However, appropriately designed non-linearity can mitigate that susceptibility while preserving the responsiveness, even leading to EP transitions [208, 209]. Bai et al. [173] designed a two-component circuit that juxtaposes the non-linearity and non-Hermitian coupling between two oscillators [Fig. 2d1]. Introducing non-linear gain results in higher-order non-linear EPs (NEPs), such as third-order NEPs, as shown in Fig. 2d2. They effectively suppress noise, even when the circuit is intentionally disturbed by a non-negligible external pulse [210]. With non-linearity, the Petermann factor can be prevented from diverging [Fig. 2d3]. Additionally, the signal-to-noise ratio reaches its lowest value at the NEP and increases as the system moves away from the NEP [Fig. 2d4].
PT symmetry-based implementations can also enhance system stability. In semiconductor technologies, fast-switching devices are widely used, and achieving both high speed and stability is highly desirable [211, 212]. Yang et al. [174] demonstrated that transient PT symmetry, as implemented in their electric circuit [Fig. 2e1], can significantly reduce unwanted oscillations during fast switching. Noise-induced perturbations in the oscillations decay more rapidly at the EP compared to in the broken or unbroken PT phases [Fig. 2e2]. For step-like dynamic switching, transient PT symmetry effectively suppresses unwanted oscillations [red in Fig. 2e3], as observed from the diminishing peak around 36 MHz in the transient PT phase [blue in Fig. 2e4] in the frequency domain.
Mathematically, EPs can also trace out interesting mathematical structures in either parameter space or real/momentum space. EPs underlie braiding in the complex eigenvalue plane [213, 214] and define knots in the band topology [215]. By designing non-Hermitian circuits involving multiple bands, Cao et al. [175] achieved higher-order EPs. The shape of knots or braids primarily depends on the winding paths around the EPs [216]. For example, a circuit with two nodes [Figs. 2f1 and f2] or three nodes [Fig. 2f3] can exhibit two-component or three-component knot topology, with double or triple twisting in the complex eigenenergy plane. Momentum-space EP knots are also described in Refs. [217, 218, 219]. In a larger parameter space, multi-band EPs can also describe mathematical singularities known as catastrophes, which have deep connections, i.e., McKay correspondence with diverse mathematical constructs such as the ADE classification of Lie algebras, finite subgroups of , platonic solids and the monodromy group of simple singularities [220, 221]. A recent circuit experiment has mapped out the swallowtail catastrophe in parameter space [222].
@@ III.1.2 Exceptional bound states as robust circuit resonances
The defectiveness of EPs causes the occupied band projector to be singular, leading to substantial non-locality in its two-point correlation functions. A prominent consequence is the appearance of special isolated eigenstates [Fig. 3a] of the real-space truncated two-point correlator matrix known as exceptional bound (EB) states [224, 223, 225, 226]. These EB states differ from topological modes, which are protected by non-trivial bulk topology, and non-Hermitian skin modes, which stem from non-trivial point gap band topology and are also boundary-localized in open-boundary systems [227, 228, 229, 230, 231, 232, 233, 234]. While most eigenvalues of the truncated two-point correlator matrix remain clustered around or , as in non-EB cases, a distinctive pair of EB states and additionally appears as long as the parent Hamiltonian is defective [Fig. 3a]. The value of scales strongly with the system size, as detailed in Refs. [224, 223]. Since the spectrum of the truncated two-point correlator matrix corresponds to fermionic occupation probabilities, EB states also lead to negative entanglement entropy in the context of free fermions [224, 223, 235], with interesting ramifications in topological [226] or competitive non-Hermitian skin effect settings [236].
@@ III.2 Non-Hermitian skin effect
One of the most intensely studied non-Hermitian phenomenon is the the non-Hermitian skin effect (NHSE), where states accumulate against spatial inhomogeneities, impurities or boundaries [237, 238, 239, 240, 241, 242, 229, 230, 227, 243, 165, 244, 245, 246, 247, 248, 249, 250, 251, 252, 253, 254, 255, 256, 257, 258, 259, 260, 261, 262, 263, 264, 265, 266, 267, 268, 269, 270, 271, 272, 273, 274, 275, 276, 277, 278, 279, 280, 281, 282, 283, 284, 285, 286, 287], challenging conventional condensed-matter insights into the correspondence between bulk and boundary properties.
In most cases, the NHSE occurs when the periodic boundary condition (PBC) spectrum encloses a finite area in the 2D complex spectral plane, which generally occurs when asymmetric couplings are present. Then, it can be shown that the open boundary condition (OBC) spectrum must assume the form of arcs or branches within the PBC spectral loop. Referring the reader to the many excellent references that elaborate on this [288, 289, 228, 231, 290, 291, 234, 241, 292, 245, 293, 294], we briefly note that the OBC spectrum, and hence the NHSE, can be understood from the Bloch PBC Hamiltonian by complex-deforming the momentum, i.e., introducing an imaginary flux [242, 295, 248, 294]:
where and represent the real and imaginary components of the momentum wavevector, respectively. The form of the complex deformation , which generically depends on , is crucial for understanding the nature of the so-called skin modes which exponentially localize against the system’s boundaries. For any Bloch wavevector with eigenvector , an appropriate complex momentum deformation would yield the OBC eigenequation
@@
Because is in general not in the Bloch (PBC) spectrum, is in general nonzero too. To determine values of which are actually allowed to exist in the OBC spectrum, we note that satisfying OBCs at both boundaries simultaneously generally requires the exact OBC eigenstate to be a superposition of two different bulk solutions with equal decay lengths . In other words, the OBC eigenenergies are those values of for which Eq. (14) simultaneously admits two different solutions with the same value of , a condition famously known as the GBZ condition [239, 296, 297, 298, 294].
Note that the GBZ condition, as well as the existence of a unique that quantifies the skin accumulation, may no longer hold when more than one NHSE channel coexists in the system. A prominent example is the critical non-Hermitian skin effect (cNHSE) [299, 300, 301, 236, 302, 303, 225, 304], where weakly coupled chains with different NHSE directions produces feedback loops that amplify certain states and lead to qualitatively different band structure descriptions at different system sizes. Appropriately designed cNHSE systems may even undergo an exceptional phase transition at a particular system size, with special scaling-induced EP defectiveness and leading to unique entanglement dips [225]. Fundamentally, the GBZ condition hinges on the assumption that the OBC eigenstate only comprise of the superposition of two bulk solutions; relaxing that generally leads to a “fragmented” GBZ in which more than one participates in the physics substantially [305, 306].
@@ III.2.1 Circuit demonstrations of one-dimensional NHSE phenomena
In a related vein, Shen et al. [308] conducted a circuit experiment involving multiple asymmetric hoppings in coupled 1D SSH chains [Fig. 4c1]. They extended this setup by connecting multiple 1D SSH chains to form a spherical geometry, paving the framework for studying the interplay between band topology and real-space lattice topology [166]. While localization behavior in various lattice geometries, including kagome and honeycomb lattices [313, 314], has been widely studied, in the spherical model [Fig. 4c2, c3], localization occurs only when a defect is introduced [pointed by red arrow in Fig. 4c4]. Additionally, Jiang et al. [315] demonstrated how circuits with multiple non-reciprocal couplings can give rise to corner modes in 2D circuits.